HIGH RESOLUTION LASER SPECTROSCOPY OF RHODIUM MONOCHLORIDE.

ALLAN G. ADAM and \underline{SCOTT A. SHEPARD}, Department of Chemistry, University of New Brunswick, Fredericton, NB, Canada E3B 6E2; WALTER J. BALFOUR and RUNHUA LI, Department of Chemistry, University of Victoria, Victoria, BC, Canada V8W 3V6.

High resolution laser spectra of rhodium monochloride (RhCl) have been acquired in the green and blue regions of the visible spectrum. The molecules were produced via laser ablation of a rhodium target rod, followed by reaction with trichloromethane in a pulsed supersonic jet. Several electronic transitions have been observed and analysed between 19,700 and 23,100 cm^{−1}. Two distinct lower state omega values have been determined: Ω = 3 which is consistent with the ^{3}Δ_{3} ground state observed for RhH and RhD ^{a}, and Ω = 2 which curiously enough seems to be the ground state of RhF ^{b}. As of yet, which Ω value represents the ground state is unknown as transitions from either of these states seem equally intense. Dispersed fluorescence (DF) scans have been taken from a number of the observed excited states yielding an estimate of 350 ± 20 cm^{−1} for a lower state vibrational frequency. The DF scans are also quite complicated indicating the presence of a number of low-lying states with energies less than 3200 cm^{−1} above the ground state. Work on the RhCl molecule is continuing and results will be discussed.

aW.J.Balfour, J. Cao, and C.X.W. Qian, J. Mol. Spectrosc. \underline{201}, 244(2000)

bA.G. Adam, W.J. Balfour, R. Li, and S.A. Shepard, to be published